### v1

# On new identities of Jacobi sums and related cyclotomic matrices

Abstract.
In this paper, using some arithmetic properties of Jacobi sums, we investigate some products involving Jacobi sums and reveal the connections between these products and certain cyclotomic matrices. In particular, as an application of our main results, we confirm a conjecture posed by Z.-W. Sun in 2019.
Key words and phrases:
Jacobi sums, cyclotomic matrices, finite fields.
2020 Mathematics Subject Classification. Primary 11L05, 15A15; Secondary 11R18, 12E20.
This research was supported by the Natural Science Foundation of China (Grant Nos. 12101321 and 12071208).
*Corresponding author.
1. Introduction
1.1. Notation
Throughout this paper, $q=p^{f}=2n+1$ is an odd prime power, where $p$ is an odd prime and $f\in\mathbb{Z}^{+}=\{1,2,\cdots\}$. The finite field with $q$ elements is denoted by $\mathbb{F}_{q}$, and let $\mathbb{F}_{q}^{\times}=\mathbb{F}_{q}\setminus\{0\}$ be the multiplicative group of all nonzero elements over $\mathbb{F}_{q}$. Also,
$\mathcal{S}_{q}=\left\{s_{0}=0,s_{1}=1,s_{2},\cdots,s_{n}\right\}=\left\{x^{2}:\ x\in\mathbb{F}_{q}\right\}$
denotes the set of all squares over $\mathbb{F}_{q}$.
Let $\widehat{\mathbb{F}_{q}^{\times}}$ denote the cyclic group of all multiplicative characters of $\mathbb{F}_{q}$. Fix a generator $\chi_{q}$ of $\widehat{\mathbb{F}_{q}^{\times}}$. Then, for any character $\chi_{q}^{k}\in\widehat{\mathbb{F}_{q}^{\times}}$, we define $\chi_{q}^{k}(0)=0$. The quadratic character $\chi_{q}^{\pm n}$ is denoted by $\phi$. Clearly,
$\phi(x)=\begin{cases}1&\mbox{if}\ x\in\mathcal{S}_{q}\setminus\{0\},\\ 0&\mbox{if}\ x=0,\\ -1&\mbox{if}\ x\in\mathbb{F}_{q}\setminus\mathcal{S}_{q}\end{cases}$
For any $\chi_{q}^{i},\chi_{q}^{j}\in\widehat{\mathbb{F}_{q}^{\times}}$, the Jacobi sum of $\chi_{q}^{i}$ and $\chi_{q}^{j}$ is defined by
$J_{q}(\chi_{q}^{i},\chi_{q}^{j}):=\sum_{x\in\mathbb{F}_{q}}\chi_{q}^{i}(x)\chi_{q}^{j}(1-x)\in\mathbb{Q}(\zeta_{q-1}),$
where $\zeta_{q-1}$ is a primitive (q1)(q-1)-th root of unity.
In addition, for any square matrix $M$ over a field $\mathbb{F}$, we use $\det M$ to denote the determinant of $M$. Also, for any integers $a<b$, the symbol $(a,b)$ denotes the set $\{a+1,\cdots,b-1\}$, and the symbol $[a,b]$ denotes the set $\{a,a+1,\cdots,b\}$.
1.2. Background and motivation
In 1827, Jacobi introduced the Jacobi sum in a letter to Gauss. Nowadays, Jacobi sums have extensive applications in number theory, and many well-known mathematicians made contributions to the theory of Jacobi sums. Readers may refer to [1] for detailed introduction on Jacobi sums.
In this paper, we mainly focus on some products involving Jacobi sums. We begin with a classical result due to Weil. Let 1<m<q11<m<q-1 be a divisor of q1q-1 with q1=lmq-1=lm. Let
$C_{m}:=\left\{[x,y,z]\in\mathbb{P}^{2}(\mathbb{F}_{q}^{{\rm alg}}):\ x^{m}+y^{m}=z^{m}\right\}$
be the projective Fermat curve defined over $\mathbb{F}_{q}$, where $\mathbb{F}_{q}^{{\rm alg}}$ is an algebraic closure of $\mathbb{F}_{q}$, and $\mathbb{P}^{2}(\mathbb{F}_{q})$ is the projective plane. Consider the zeta function of $C_{n}$ defined by
$\zeta_{C_{m}}(t):=\exp\left(\sum_{r\in\mathbb{Z}^{+}}\frac{N(q^{r})\cdot t^{r}}{r}\right),$
where $N(q^{r})$ is the number of $\mathbb{F}_{q^{r}}$-rational points on $C_{m}$. As $C_{m}$ is a nonsingular absolutely irreducible curve of genus (m1)(m2)/2(m-1)(m-2)/2, by the Weil theorem, there is an integral polynomial $P_{m}(t)$ with $\deg(P_{m}(t))=(m-1)(m-2)$ such that
$\zeta_{C_{m}}=\zeta_{C_{m}}(t)=\frac{P_{m}(t)}{(1-t)(1-qt)}.$
Weil proved that the integral polynomial $P_{m}(t)$ has close relations with products of Jacobi sums, that is,
$P_{m}(t)=\prod_{\begin{subarray}{c}i,j\in[1,m-1]\\ i+j\not\equiv 0\ ({\rm{mod}}\ m)\end{subarray}}\left(1+J_{q}(\chi_{q}^{li},\chi_{q}^{lj})t\right)\in\mathbb{Z}[t].$
We next turn to Carlitz’s result. Let $(\frac{\cdot}{p})$ be the Legendre symbol, where $p$ is an odd prime. Carlitz [2] proved that the determinant of the matrix
(1.1) $C_{p}:=\left[\left(\frac{i-j}{p}\right)\right]_{1\leq i,j\leq p-1}$
can be represented by a product of Jacobi sums, that is,
$\det C_{p}=(-1)^{\frac{p-1}{2}}\cdot\prod_{k=1}^{p-1}J_{p}(\phi,\chi_{p}^{k})=p^{\frac{p-3}{2}}.$
Along this line, recently, the first author and Wang [14, Theorem 1.1] showed that if $q\equiv 3\ ({\rm{mod}}\ 4)$ is a prime power, then
$\det\left[\chi_{q}^{r}(s_{i}+s_{j})+\chi_{q}^{r}(s_{i}-s_{j})\right]_{1\leq i,j\leq(q-1)/2}=\prod_{k=0}^{(q-3)/2}J_{q}(\chi_{q}^{r},\chi_{q}^{2k})$
for any $r\in[1,q-2]$.
It might be worth mentioning here that the above two results fully illustrate that the explicit values of certain products of Jacobi sums can be represented by the determinants of some cyclotomic matrices.
We next introduce some results due to Z.-W. Sun. Inspired by Carlitz’s matrix $C_{p}$ defined by (1.1), Z.-W. Sun [11] studied the matrix
$S_{p}:=\left[\left(\frac{i^{2}+j^{2}}{p}\right)\right]_{1\leq i,j\leq(p-1)/2},$
and conjectured that $\det S_{p}$ is a nonzero square of some integer whenever the prime $p\equiv 3\ ({\rm{mod}}\ 4)$. This conjecture was later confirmed by Alekseyev and Krachun. When $p\equiv 1\ ({\rm{mod}}\ 4)$, if we write $p=a^{2}+4b^{2}$ with $a,b\in\mathbb{Z}$ and $a\equiv 1\ ({\rm{mod}}\ 4)$, then Cohen, Sun and Vsemirnov conjectured that $\det S_{p}/a$ is a nonzero square. The first author [13] proved this conjecture and revealed the connections between $\det S_{p}$ and the product
$R_{p}(\chi_{p}):=\prod_{k\in(0,(p-1)/4)}\left(J_{p}(\phi,\chi_{p}^{k})+J_{p}(\phi,\chi_{p}^{-k})\right),$
which concerns the real part of the Jacobi sum $J_{p}(\phi,\chi_{p}^{k})$.
Motivated by the above results, in this paper, we mainly investigate the product
(1.2) $I_{q}(\chi_{q}):=\prod_{k\in(0,n/2)}\left(J_{q}(\phi,\chi_{q}^{k})-J_{q}(\phi,\chi_{q}^{-k})\right),$
which involves the imaginary part of the Jacobi sum $J_{q}(\phi,\chi_{q}^{k})$.
1.3. Main results
Now we state our two theorems concerning $I_{q}(\chi_{q})$. Suppose first $q\equiv 3\ ({\rm{mod}}\ 4)$ with $q=2n+1$. Note that the matrix
$N_{q}(-1):=\left[\phi(s_{i}-s_{j})\right]_{2\leq i,j\leq n}$
is a skew-symmetric matrix of even order with integer entries. Then by Stembridge’s result [10, Proposition 2.2], $\det N_{q}(-1)$ is indeed a square of some integer. Our first theorem reveals the relationship between $I_{q}(\chi_{q})$ and $\det N_{q}(-1)$.
Theorem 1.1.
Let $q=2n+1$ be an odd prime power with $q\equiv 3\ ({\rm{mod}}\ 4)$, and let $\chi_{q}$ be a generator of $\widehat{\mathbb{F}_{q}^{\times}}$. Then
$x_{q}:=\frac{I_{q}(\chi_{q})}{\sqrt{(-1)^{\frac{n-1}{2}}\cdot n}}\in\mathbb{Z}.$
Moreover, we have
$x_{q}^{2}=2^{n-1}\cdot\det\left[\phi(s_{i}-s_{j})\right]_{2\leq i,j\leq n}.$
We next turn to the case $q\equiv 1\ ({\rm{mod}}\ 4)$. Inspired by Theorem 1.1, when $q\equiv 1\ ({\rm{mod}}\ 4)$, one may expect that $I_{q}(\chi_{q})$ has a close relationship with certain cyclotomic matrix involving $\phi$. In fact, we will see that $I_{q}(\chi_{q})$ is not only related to some cyclotomic matrix, but also to the number of $\mathbb{F}_{q}$-rational points on certain elliptic curve.
For any element $d\in\mathbb{F}_{q}^{\times}\setminus\mathcal{S}_{q}$, define the matrix
$T_{q}(d):=\left[\phi(s_{i}+ds_{j})\right]_{0\leq i,j\leq n}.$
Also, let $X_{d}$ be the elliptic curve defined by the equation $y^{2}=dx^{3}+x$ over $\mathbb{F}_{q}$, and let
$X_{d}(\mathbb{F}_{q})=\left\{(x,y)\in\mathbb{F}_{q}\times\mathbb{F}_{q}:\ y^{2}=dx^{3}+x\right\}\cup\left\{\infty\right\}$
be the set of all $\mathbb{F}_{q}$-rational points on $X_{d}$. Then it is known that
$X_{d}(\mathbb{F}_{q})=q+1-a_{d}(q),$
where #S\#S denotes the cardinality of a finite set $S$, and
(1.3) $a_{d}(q)=-\sum_{x\in\mathbb{F}_{q}}\phi(dx^{3}+x)$
is the trace of Frobenius map.
Let $\bm{i}$} be the primitive $4$-th root of unity with argument $\pi/2$. Now we state our second theorem.
Theorem 1.2.
Let $q=2n+1\geq 7$ be an odd prime power with $q\equiv 1\ ({\rm{mod}}\ 4)$, and $\chi_{q}$ be a generator of $\widehat{\mathbb{F}_{q}^{\times}}$. Then
$y_{q}:={\bm{i}}^{\frac{n-2}{2}}\cdot I_{q}(\chi_{q})\cdot\sqrt{q-1}\in\mathbb{Z}.$
Moreover, for any $d\in\mathbb{F}_{q}^{\times}\setminus\mathcal{S}_{q}$, we have
$a_{d}(q)\cdot y_{q}^{2}=2^{n}\cdot\det T_{q}(d).$
In 2019, Z.-W. Sun [11, Conjecture 4.2(ii)] posed the following conjecture.
Conjecture 1.1.
Let $p\equiv 1\ ({\rm{mod}}\ 4)$ be a prime with $p=c_{p}^{2}+4b_{p}^{2}$, where $c_{p}$ and $b_{p}$ are positive integers. Then for any $d\in\mathbb{F}_{p}^{\times}\setminus\mathcal{S}_{p}$, the number $|\det T_{p}(d)|/(2^{(p-1)/2}b_{p})$ is a positive integral square not depending on $d$.
As an application of Theorem 1.2, we can confirm this conjecture.
Corollary 1.1.
Conjecture 1.1 holds.
1.4. Outline of this paper
In Section 2, we will prove some necessary lemmas. In particular, we will prove an important lemma, which can be viewed as a variant of the well-known Gauss lemma on quadratic residues. Our main results will be proved in Section 3 and Section 4 respectively. In Section 5, we shall briefly discuss a recent conjecture posed by Z.-W. Sun, which can be viewed as a refinement of Conjecture 1.1.
2. Preparations for the proofs of main results
2.1. On a variant of the Gauss lemma
We first introduce some notations. For any real number $y$, let
$\lfloor y\rfloor:=\max\{u\in\mathbb{Z}:\ u\leq y\}.$
For any positive integer $m$ and any integer $x$, let $\{x\}_{m}$ denote the unique integer $r\in[0,m-1]$ with $r\equiv x\ ({\rm{mod}}\ m)$.
Let $p$ be an odd prime and $a\in\mathbb{Z}$ with $\gcd(a,p)=1$. Then the famous Gauss lemma (cf. [4, p. 52]) says that
$\left(\frac{a}{p}\right)=(-1)^{\#\left\{1\leq x\leq\frac{p-1}{2}:\ \{ax\}_{p}>\frac{p}{2}\right\}}.$
This result was later extended to Jacobi symbol by Jenkins. Let $s$ be a positive odd integer and $a\in\mathbb{Z}$ with $\gcd(a,s)=1$. Then Jenkins [5] proved that
(2.1) $\left(\frac{a}{s}\right)=(-1)^{\#\left\{1\leq x\leq\frac{s-1}{2}:\ \{ax\}_{s}>\frac{s}{2}\right\}},$
where $(\frac{\cdot}{s})$ is the Jacobi symbol. Using this, it is easy to verify the following lemma.
Lemma 2.1.
Let $s$ be a positive odd integer and $a\in\mathbb{Z}$ with $\gcd(a,s)=1$. Then
$(-1)^{K}=\left(\frac{a}{s}\right),$
where
$K=\sum_{x=1}^{(s-1)/2}\left\lfloor\frac{2ax}{s}\right\rfloor.$
The second author [8, Lemma 1] obtained the following result.
Lemma 2.2.
Let $m\geq 2$ be an integer and $a\in\mathbb{Z}$ with $\gcd(a,m)=1$. Suppose that $1\leq i<j\leq m-1$ and $\{ai\}_{m}>\{aj\}_{m}$. Then
$\left\lfloor\frac{aj}{m}\right\rfloor-\left\lfloor\frac{ai}{m}\right\rfloor-\left\lfloor\frac{a(j-i)}{m}\right\rfloor=1.$
We also need the following result.
Lemma 2.3.
Let $m\geq 2$ be an even integer and $a$ be a positive integer with $\gcd(a,m)=1$. Then
$\sum_{x=1}^{(a-1)/2}\left\lfloor\frac{2mx}{a}\right\rfloor+\sum_{y=1}^{m}\left\lfloor\frac{ay}{2m}\right\rfloor=\frac{(a-1)m}{2}.$
Proof.
By symmetry, one can verify that the sum
$\sum_{x=1}^{(a-1)/2}\left\lfloor\frac{2mx}{a}\right\rfloor+\sum_{y=1}^{m}\left\lfloor\frac{ay}{2m}\right\rfloor$
exactly counts the number of integral points of the set
$\left\{(x,y)\in\mathbb{R}\times\mathbb{R}:\ 0<x<a/2\ \text{and}\ 1\leq y\leq m\right\}.$
By this, our lemma holds immediately. ∎
Now we are in a position to prove the following variant of Gauss lemma, which will play an important role in the proof of main theorems.
Lemma 2.4.
Let $m\geq 2$ be an even integer and $a$ be a positive integer with $\gcd(a,m)=1$. Then
$(-1)^{\frac{(a-1)(m-2)}{4}+\#\left\{1\leq x\leq\frac{m-2}{2}:\ \{ax\}_{2m}>m\right\}}=\left(\frac{2m}{a}\right).$
Proof.
Note that $\{am\}_{2m}=m$. By Lemma 2.2, one can verify that
$\#\left\{1\leq x\leq\frac{m-2}{2}:\ \{ax\}_{2m}>m\right\}$
$=$ $\sum_{x=1}^{(m-2)/2}\left(\left\lfloor\frac{am}{2m}\right\rfloor-\left\lfloor\frac{ax}{2m}\right\rfloor-\left\lfloor\frac{a(m-x)}{2m}\right\rfloor\right)$
(2.2) $=$ $\frac{(a-1)(m-2)}{4}+\left\lfloor\frac{a}{4}\right\rfloor-\sum_{y=1}^{m-1}\left\lfloor\frac{ay}{2m}\right\rfloor.$
Meanwhile, by Lemma 2.3
(2.3) $\sum_{x=1}^{(a-1)/2}\left\lfloor\frac{2mx}{a}\right\rfloor+\sum_{y=1}^{m-1}\left\lfloor\frac{ay}{2m}\right\rfloor=\frac{(a-1)(m-1)}{2}.$
Suppose first $a\equiv 1\ ({\rm{mod}}\ 4)$. Combining (2.1) and (2.3) with Lemma 2.1, we obtain
$(-1)^{\#\left\{1\leq x\leq\frac{m-2}{2}:\ \{ax\}_{2m}>m\right\}}$ $=(-1)^{\frac{a-1}{4}+\sum_{y=1}^{m-1}\left\lfloor\frac{ay}{2m}\right\rfloor}$
$=(-1)^{\frac{a-1}{4}+\sum_{x=1}^{(a-1)/2}\left\lfloor\frac{2mx}{a}\right\rfloor}$
$=\left(\frac{2m}{a}\right).$
Now consider the case $a\equiv 3\ ({\rm{mod}}\ 4)$. By (2.1), (2.3) and Lemma 2.1 again
$(-1)^{\#\left\{1\leq x\leq\frac{m-2}{2}:\ \{ax\}_{2m}>m\right\}}$ $=(-1)^{\frac{m-2}{2}+\frac{a-3}{4}+\sum_{y=1}^{m-1}\left\lfloor\frac{ay}{2m}\right\rfloor}$
$=(-1)^{\frac{m-2}{2}+\frac{a-3}{4}+1+\sum_{x=1}^{(a-1)/2}\left\lfloor\frac{2mx}{a}\right\rfloor}$
$=(-1)^{\frac{m-2}{2}}\cdot\left(\frac{2m}{a}\right).$
In view of the above, we have completed the proof. ∎
2.2. Some known results on quadratic Gauss sums over $\mathbb{Z}/m\mathbb{Z}$
Let $p>2$ be an odd prime. Gauss first determined the explicit value of the quadratic Gauss sum over $\mathbb{Z}/p\mathbb{Z}$, that is,
$\sum_{x\in\mathbb{Z}/p\mathbb{Z}}\left(\frac{x}{p}\right)e^{2\pi{\bm{i}}x/p}=\sum_{x\in\mathbb{Z}/p\mathbb{Z}}e^{2\pi{\bm{i}}x^{2}/p}=\sqrt{(-1)^{\frac{p-1}{2}}p}.$
The above result can be extended to the quadratic Gauss sum over $\mathbb{Z}/m\mathbb{Z}$, where $m$ is an arbitrary positive integer (cf. [7, pp. 85-87]).
Lemma 2.5.
Let $m$ be a positive integer. Then
$\sum_{x\in\mathbb{Z}/m\mathbb{Z}}e^{2\pi{\bm{i}}x^{2}/m}=\begin{cases}0&\mbox{if}\ m\equiv 2\ ({\rm{mod}}\ 4),\\ (1+{\bm{i}})\sqrt{m}&\mbox{if}\ m\equiv 0\ ({\rm{mod}}\ 4),\\ \sqrt{(-1)^{\frac{m-1}{2}}m}&\mbox{if}\ m\equiv 1\ ({\rm{mod}}\ 2).\end{cases}$
Moreover, if $m$ is a positive odd integer, then
$\sum_{x\in\mathbb{Z}/m\mathbb{Z}}e^{2\pi{\bm{i}}ax^{2}/m}=\left(\frac{a}{m}\right)\cdot\sum_{x\in\mathbb{Z}/m\mathbb{Z}}e^{2\pi{\bm{i}}x^{2}/m}$
for any $a\in\mathbb{Z}$ with $\gcd(a,m)=1$.
To state our next result, we introduce the Kronecker symbol, which is indeed an extension of Jacobi symbol. In fact, the Kronecker symbol is a function
$\left(\frac{\bullet}{\bullet}\right):\ \left(\mathbb{Z}\setminus\{0\}\right)\times\left(\mathbb{Z}\setminus\{0\}\right)\longrightarrow\left\{-1,0,1\right\}$
defined by the following conditions:
(i) $\left(\frac{b}{a_{1}}\right)\left(\frac{b}{a_{2}}\right)=\left(\frac{b}{a_{1}a_{2}}\right)$ for any nonzero integers $a_{1},a_{2},b$;
(ii) If $a$ is a positive odd integer, then $(\frac{b}{a})$ coincides with the Jacobi symbol;
(iii) For any nonzero integer $b$, we have $(\frac{b}{-1})=\mathrm{sign}(b)=b/|b|$;
(iv) For any nonzero integer $b$, we have
$\left(\frac{b}{2}\right)=\begin{cases}0&\mbox{if}\ b\equiv 0\ ({\rm{mod}}\ 2),\\ (\frac{2}{b})&\mbox{if}\ b\equiv 1\ ({\rm{mod}}\ 2).\end{cases}$
Now we state the following known result on character sums related to the Kronecker symbol (cf. [3, Corollary 2.1.47]).
Lemma 2.6.
Let $K$ be a quadratic number field with discriminant $D$ and let $m=|D|$. Then
$\sqrt{D}=\sum_{x\in(\mathbb{Z}/m\mathbb{Z})\setminus\{0\ {\rm mod}\ m\mathbb{Z}\}}\left(\frac{D}{x}\right)e^{2\pi{\bm{i}}x/m}.$
2.3. Circulant and almost circulant matrices
Let $m\geq 2$ be an integer and let
${\bm{v}}=(c_{0},c_{1},\cdots,c_{m-1})\in\mathbb{C}^{m}.$
Then the circulant matrix $C({\bm{v}})$ of the vector $\bm{v}$} is an $m\times m$ matrix defined by
$C({\bm{v}}):=\left[c_{j-i}\right]_{0\leq i,j\leq m-1},$
where $c_{x}=c_{\{x\}_{m}}$ for any integer $x$. More precisely,
$C({\bm{v}})=\left[\begin{array}[]{ccccc}c_{0}&c_{1}&\cdots&c_{m-2}&c_{m-1}\\ c_{m-1}&c_{0}&\cdots&c_{m-3}&c_{m-2}\\ \vdots&\vdots&\ddots&\vdots&\vdots\\ c_{2}&c_{3}&\cdots&c_{0}&c_{1}\\ c_{1}&c_{2}&\cdots&c_{m-1}&c_{0}\end{array}\right].$
Circulant matrices have extensively applications in number theory, combinatorics, as well as coding theory. Readers may refer to the survey paper [6] for more details on circulant matrices.
In 2025, the first author and Wang investigated the matrix
$W({\bm{v}}):=\left[c_{j-i}\right]_{1\leq i,j\leq m-1},$
which is obtained by deleting the first row and the first column of $C({\bm{v}})$. The matrix $W({\bm{v}})$ is called the almost circulant matrix of $\bm{v}$}. For $W({\bm{v}})$, the first author and Wang [14, Theorem 4.1] obtained the following result.
Lemma 2.7.
Let notations be as above and let $\lambda_{1},\lambda_{2},\cdots,\lambda_{m}$ be exactly all the eigenvalues of the circulant matrix $C({\bm{v}})$. Then
$\det W({\bm{v}})=\frac{1}{m}\sum_{l=1}^{m}\prod_{k\in[1,m]\setminus\{l\}}\lambda_{k}.$
In particular, if $\lambda_{m}=0$, then
$\det W({\bm{v}})=\frac{1}{m}\lambda_{1}\lambda_{2}\cdots\lambda_{m-1}.$
2.4. Some necessary lemmas
Here we introduce some lemmas, which will be used frequently in the proofs of main theorems. Recall that $q=2n+1$ is an odd prime power and $\chi_{q}$ is a generator of $\widehat{\mathbb{F}_{q}^{\times}}$. We begin with the following result.
Lemma 2.8.
Let notations be as above. Then
$J_{q}(\chi_{q}^{i},\chi_{q}^{j})=(-1)^{i}\cdot J_{q}(\chi_{q}^{i},\chi_{q}^{-(i+j)}),$
for any integers $i$ and $j$. In particular,
$J_{q}(\phi,\chi_{q}^{j})=\phi(-1)\cdot J_{q}(\phi,\chi_{q}^{n-j}).$
Proof.
By the definition of Jacobi sums, one can verify that
$(-1)^{i}\cdot J_{q}(\chi_{q}^{i},\chi_{q}^{-(i+j)})$ $=\sum_{x\in\mathbb{F}_{q}^{\times}}\chi_{q}^{i}(x-1)\chi_{q}^{i+j}\left(\frac{1}{x}\right)$
$=\sum_{x\in\mathbb{F}_{q}^{\times}}\chi_{q}^{i}\left(1-\frac{1}{x}\right)\chi_{q}^{j}\left(\frac{1}{x}\right)$
$=\sum_{x\in\mathbb{F}_{q}^{\times}}\chi_{q}^{i}(1-x)\chi_{q}^{j}(x)$
$=J_{q}(\chi_{q}^{i},\chi_{q}^{j}).$
This completes the proof. ∎
Recall that
$\mathcal{S}_{q}=\left\{s_{0}=0,s_{1}=1,s_{2},\cdots,s_{n}\right\}=\left\{x^{2}:\ x\in\mathbb{F}_{q}\right\}.$
Throughout out the remaining part of this paper, for $k=1,2,\cdots,n$, we define
$\lambda_{k}(d):=\sum_{s\in\mathcal{S}_{q}}\phi(1+ds)\chi_{q}^{k}(s).$
Using these notations, we have the following result.
Lemma 2.9.
Let $d\in\mathbb{F}_{q}^{\times}$ and let the matrix
$M_{q}(d):=\left[\phi(s_{i}+ds_{j})\right]_{1\leq i,j\leq n}.$
Then the following results hold.
(i) $\lambda_{1}(d),\lambda_{2}(d),\cdots,\lambda_{n}(d)$ are exactly all the eigenvalues of $M_{q}(d)$.
(ii) Moreover, if $d\in\mathbb{F}_{q}^{\times}\setminus\mathcal{S}_{q}$, then for $k=1,2,\cdots,n$, we have
$\chi_{q}^{k}(d)\cdot\lambda_{k}(d)=\frac{(-1)^{k}}{2}\left(J_{q}(\phi,\chi_{q}^{k})-J_{q}(\phi,\chi_{q}^{-k})\right).$
Proof.
(i) For any $k\in[1,n]$, define the column vector
${\bm{v}}_{k}:=\left(\chi_{q}^{k}(s_{1}),\chi_{q}^{k}(s_{2}),\cdots,\chi_{q}^{k}(s_{n})\right)^{T}.$
Then one can verify that
$\sum_{j\in[1,n]}\phi(s_{i}+ds_{j})\chi_{q}^{k}(s_{j})$ $=\sum_{j\in[1,n]}\phi(1+ds_{j}/s_{i})\chi_{q}^{k}(s_{j}/s_{i})\chi_{q}^{k}(s_{i})$
$=\sum_{j\in[1,n]}\phi(1+ds_{j})\chi_{q}^{k}(s_{j})\cdot\chi_{q}^{k}(s_{i})$
$=\lambda_{k}(d)\cdot\chi_{q}^{k}(s_{i}).$
This implies
$M_{q}(d){\bm{v}}_{k}=\lambda_{k}(d){\bm{v}}_{k}$
for any $k\in[1,n]$. As vectors ${\bm{v}}_{1},{\bm{v}}_{2},\cdots,{\bm{v}}_{n}$ are linearly independent over $\mathbb{C}$, the numbers $\lambda_{1}(d),\lambda_{2}(d),\cdots,\lambda_{n}(d)$ are exactly all the eigenvalues of $M_{q}(d)$. Hence (i) holds.
(ii) Suppose now $d\in\mathbb{F}_{q}^{\times}\setminus\mathcal{S}_{q}$, i.e., $\phi(d)=-1$. Then the conjugate of the complex number $\chi_{q}^{k}(d)\cdot\lambda_{k}(d)$ is equal to
$\overline{\chi_{q}^{k}(d)\cdot\lambda_{k}(d)}$ $=\sum_{s\in\mathcal{S}_{q}}\phi(1+ds)\chi_{q}^{-k}(ds)$
$=-\chi_{q}^{k}(d)\sum_{s\in\mathcal{S}_{q}}\phi(d+d^{2}s)\chi_{q}^{-k}(d^{2}s)$
$=-\chi_{q}^{k}(d)\sum_{s\in\mathcal{S}_{q}\setminus\{0\}}\phi\left(d+\frac{1}{s}\right)\chi_{q}^{k}(s)$
$=-\chi_{q}^{k}(d)\sum_{s\in\mathcal{S}_{q}\setminus\{0\}}\phi(1+ds)\chi_{q}^{k}(s)$
$=-\chi_{q}^{k}(d)\cdot\lambda_{k}(d).$
This implies that
$\chi_{q}^{k}(d)\cdot\lambda_{k}\in\mathbb{R}{\bm{i}}:=\left\{x{\bm{i}}:\ x\in\mathbb{R}\right\},$
i.e., $\chi_{q}^{k}(d)\cdot\lambda_{k}$ is either 0 or a purely imaginary number. On the other hand, by (i), the numbers $\lambda_{1}(1),\lambda_{2}(1),\cdots,\lambda_{n}(1)$ are precisely all the eigenvalues of the real symmetric matrix $M_{q}(1)$. Thus, $\lambda_{k}(1)\in\mathbb{R}$ for any $k\in[1,n]$.
As $d\in\mathbb{F}_{q}^{\times}\setminus\mathcal{S}_{q}$, we clearly have
$\left\{ds:\ s\in\mathcal{S}_{q}\setminus\{0\}\right\}\cap\mathcal{S}_{q}=\emptyset,$
and
$\left\{ds:\ s\in\mathcal{S}_{q}\setminus\{0\}\right\}\cup\mathcal{S}_{q}=\mathbb{F}_{q}.$
Using this, we obtain
$\chi_{q}^{k}(d)\cdot\lambda_{k}(d)+\lambda_{k}(1)$ $=\sum_{s\in\mathcal{S}_{q}}\phi(1+ds)\chi_{q}^{k}(ds)+\sum_{s\in\mathcal{S}_{q}}\phi(1+s)\chi_{q}^{k}(s)$
$=\sum_{x\in\mathbb{F}_{q}}\phi(1+x)\chi_{q}^{k}(x)$
$=(-1)^{k}\cdot J_{q}(\phi,\chi_{q}^{k}).$
Noting that $\chi_{q}^{k}(d)\cdot\lambda_{k}(d)\in\mathbb{R}{\bm{i}}$ and $\lambda_{k}(1)\in\mathbb{R}$, we obtain
(2.4) $\chi_{q}^{k}(d)\cdot\lambda_{k}(d)=\frac{(-1)^{k}}{2}\left(J_{q}(\phi,\chi_{q}^{k})-J_{q}(\phi,\chi_{q}^{-k})\right)$
for any $k\in[1,n]$.
In view of the above, we have completed the proof. ∎
For any $d\in\mathbb{F}_{q}^{\times}\setminus\mathcal{S}_{q}$, recall that the matrix
$T_{q}(d)=\left[\phi(s_{i}+ds_{j})\right]_{0\leq i,j\leq n}.$
We conclude this section with the following result.
Lemma 2.10.
Let $q=2n+1\equiv 1\ ({\rm{mod}}\ 4)$ be an odd prime power. Then for any $d\in\mathbb{F}_{q}^{\times}\setminus\mathcal{S}_{q}$, the $n+1$ numbers $\pm\sqrt{-n},\lambda_{1}(d),\lambda_{2}(d),\cdots,\lambda_{n-1}(d)$ are exactly all the eigenvalues of $T_{q}(d)$.
Proof.
For $k=-1,0,1,\cdots,n-1$, the column vector ${\bm{u}}_{k}$ is defined by
$\begin{cases}{\bm{u}}_{k}=\left((-1)^{k+1}\cdot\sqrt{-n},1,1,\cdots,1\right)^{T}&\mbox{if}\ k=-1,0,\\ {\bm{u}}_{k}=\left(0,\chi_{q}^{k}(s_{1}),\chi_{q}^{k}(s_{2}),\cdots,\chi_{q}^{k}(s_{n})\right)^{T}&\mbox{if}\ k\in[1,n-1].\end{cases}$
It is easy to see that the vectors ${\bm{u}}_{-1},{\bm{u}}_{0},{\bm{u}}_{1},\cdots,{\bm{u}}_{n-1}$ are linearly independent over $\mathbb{C}$.
For k=1,0k=-1,0, we clearly have
$T_{q}(d){\bm{u}}_{k}=(-1)^{k+1}\sqrt{-n}{\bm{u}}_{k}.$
Also, with essentially the same method appeared in the proof of Lemma 2.9, we obtain
$T_{q}(d){\bm{u}}_{k}=\lambda_{k}(d){\bm{u}}_{k}$
for any $k\in[1,n-1]$. In view of the above, $\pm\sqrt{-n},\lambda_{1}(d),\lambda_{2}(d),\cdots,\lambda_{n-1}(d)$ are precisely all the eigenvalues of $T_{q}(d)$. This completes the proof. ∎
3. Proofs of Theorem 1.1
Proof of Theorem 1.1. Let $q\equiv 3\ ({\rm{mod}}\ 4)$ be an odd prime power with n=(q1)/2n=(q-1)/2. We first prove that
$x_{q}=\frac{I_{q}(\chi_{q})}{\sqrt{(-1)^{\frac{n-1}{2}}\cdot n}}\in\mathbb{Q}.$
Let $\sigma_{a}\in{\rm Gal}(\mathbb{Q}(\zeta_{q-1})/\mathbb{Q})$ be an arbitrary automorphism with $a\in\mathbb{Z}$, $\gcd(a,q-1)=1$ and $\sigma_{a}(\zeta_{q-1})=\zeta_{q-1}^{a}$. For $j=1,2,3,4$, define the set
$S_{j}(a):=\left\{k\in(0,n/2):\ \frac{j-1}{2}n<\{ka\}_{2n}<\frac{j}{2}n\right\}.$
As $q\equiv 3\ ({\rm{mod}}\ 4)$, applying Lemma 2.8, we obtain
$J_{q}(\phi,\chi_{q}^{k})=-1\cdot J_{q}(\phi,\chi_{q}^{\pm n-k}).$
By this, one can verify that
(3.1) $\sigma_{a}\left(I_{q}(\chi_{q})\right)=(-1)^{\#S_{2}(a)\cup S_{4}(a)}\cdot I_{q}(\chi_{q}).$
Note that
$k_{1}a\not\equiv k_{2}a\ ({\rm{mod}}\ n)$
for any $k_{1},k_{2}\in(0,n/2)$ with $k_{1}\neq k_{2}$. Thus,
$S_{2}(a)\cup S_{4}(a)=\left\{k\in(0,n/2):\ \frac{n}{2}<\{ka\}_{n}<n\right\}.$
Combining this with (2.1) and (3.1), we immediately obtain
(3.2) $\sigma_{a}\left(I_{q}(\chi_{q})\right)=\left(\frac{a}{n}\right)\cdot I_{q}(\chi_{q}).$
On the other hand, by Lemma 2.5
(3.3) $\sigma_{a}\left(\sqrt{(-1)^{\frac{n-1}{2}}n}\right)=\left(\frac{a}{n}\right)\cdot\sqrt{(-1)^{\frac{n-1}{2}}n}.$
By (3.2) and (3.3), we see that
$\sigma_{a}(x_{q})=x_{q}$
for any $\sigma_{a}\in{\rm Gal}(\mathbb{Q}(\zeta_{q-1})/\mathbb{Q})$. Hence $x_{q}\in\mathbb{Q}$ by the Galois theory.
We next evaluate $x_{q}^{2}$. Fix a generator $g$ of cyclic group $\mathbb{F}_{q}^{\times}$ and let $s_{j}=g^{2j}$ for $j=0,1,\cdots,n-1$. Then the matrix
$M_{q}(-1)=\left[\phi(s_{i}-s_{j})\right]_{1\leq i,j\leq n}=\left[\phi(1-g^{2j-2i})\right]_{0\leq i,j\leq n-1}$
is indeed a circulant matrix $C({\bm{v}})$ of the vector
${\bm{v}}=\left(\phi(1-s_{1}),\phi(1-s_{2}),\cdots,\phi(1-s_{n})\right).$
Hence
$N_{q}(-1)=\left[\phi(s_{i}-s_{j})\right]_{2\leq i,j\leq n}=\left[\phi(1-g^{2j-2i})\right]_{1\leq i,j\leq n-1}$
is the almost circulant matrix of $\bm{v}$}. As $q\equiv 3\ ({\rm{mod}}\ 4)$, we have $1\not\in\mathcal{S}_{q}$. Note that $\lambda_{n}(-1)=0$. Applying Lemma 2.9 and Lemma 2.7, we obtain
$\det N_{q}(-1)=\frac{1}{n}\prod_{k\in[1,n-1]}\lambda_{k}(-1)=\frac{1}{2^{n-1}\cdot n}\prod_{k\in[1,n-1]}\left(J_{q}(\phi,\chi_{q}^{k})-J_{q}(\phi,\chi_{q}^{-k})\right).$
Using Lemma 2.8 again, the above equality implies
(3.4) $2^{n-1}\cdot\det N_{q}(-1)=x_{q}^{2}\in\mathbb{Z}.$
As $x_{q}\in\mathbb{Q}$, by (3.4) we see that $x_{q}\in\mathbb{Z}$.
In view of the above, we have completed the proof.∎
4. Proofs of Theorem 1.2 and its corollary
Proof of Theorem 1.2. Let $q=2n+1\geq 7$ be an odd prime power with $q\equiv 1\ ({\rm{mod}}\ 4)$. We first show that
$y_{q}={\bm{i}}^{\frac{n-2}{2}}\cdot I_{q}(\chi_{q})\cdot\sqrt{q-1}\in\mathbb{Z}.$
Let $\sigma_{a}\in{\rm Gal}(\mathbb{Q}(\zeta_{q-1})/\mathbb{Q})$ be an arbitrary automorphism with $a\in[1,q-1]$, $\gcd(a,q-1)=1$ and $\sigma_{a}(\zeta_{q-1})=\zeta_{q-1}^{a}$. For any $j\in[1,4]$, let
$S_{j}(a):=\left\{k\in(0,n/2):\ \frac{j-1}{2}n<\{ka\}_{2n}<\frac{j}{2}n\right\}.$
By Lemma 2.8 and $\phi(-1)=1$, we have
$J_{q}(\phi,\chi_{q}^{k})=J_{q}(\phi,\chi_{q}^{n-k}).$
By this and Lemma 2.4, one can verify that
(4.1) $\sigma_{a}\left({\bm{i}}^{\frac{n-2}{2}}\cdot I_{q}(\chi_{q})\right)=(-1)^{\frac{(a-1)(n-2)}{4}+\#S_{3}(a)\cup S_{4}(a)}\cdot{\bm{i}}^{\frac{n-2}{2}}\cdot I_{q}(\chi_{q})=\left(\frac{q-1}{a}\right)\cdot{\bm{i}}^{\frac{n-2}{2}}\cdot I_{q}(\chi_{q}).$
On the other hand, by Lemma 2.6 we see that
(4.2) $\sigma_{a}\left(\sqrt{q-1}\right)=\left(\frac{q-1}{a}\right)\cdot\sqrt{q-1}.$
By (4.1) and (4.2), for any $\sigma_{a}\in{\rm Gal}(\mathbb{Q}(\zeta_{q-1})/\mathbb{Q})$ we have
$\sigma_{a}(y_{q})=y_{q},$
and hence $y_{q}\in\mathbb{Q}$ by the Galois theory. As $y_{q}$ is an algebraic integer, we further obtain $y_{q}\in\mathbb{Z}$.
We next consider $y_{q}^{2}$. Recall that the matrix
$T_{d}(q)=\left[\phi(s_{i}+ds_{j})\right]_{0\leq i,j\leq n}.$
By Lemma 2.8, Lemma 2.9 and Lemma 2.10, one can verify that
$\det T_{q}(d)$ $=n\cdot\lambda_{n/2}(d)\cdot\prod_{k\in[1,n-1]\setminus\{n/2\}}\lambda_{k}(d)$
$=n\cdot\lambda_{n/2}(d)\cdot\prod_{k\in[1,n-1]\setminus\{n/2\}}\frac{(-1)^{k}\chi_{q}^{-k}(d)}{2}\left(J_{q}(\phi,\chi_{q}^{k})-J_{q}(\phi,\chi_{q}^{-k})\right)$
$=\frac{n\cdot\lambda_{n/2}(d)}{2^{n-2}}\cdot(-1)^{\frac{n-2}{2}}\cdot I_{q}(\chi_{q})^{2}$
(4.3) $=\frac{\lambda_{n/2}(d)}{2^{n-1}}\cdot y_{q}^{2}.$
Note also that
$2\cdot\lambda_{n/2}(d)$ $=2\sum_{j\in[1,n]}\phi(1+ds_{j})\chi_{q}^{n/2}(s_{j})$
$=2\sum_{j\in[1,n]}\phi(1+ds_{j})\phi\left(\sqrt{s_{j}}\right)$
$=\sum_{j\in[1,n]}\phi(1+ds_{j})\phi\left(\sqrt{s_{j}}\right)+\sum_{j\in[1,n]}\phi(1+ds_{j})\phi\left(-\sqrt{s_{j}}\right)$
$=\sum_{x\in\mathbb{F}_{q}}\phi(1+dx^{2})\phi(x),$
where $\sqrt{s_{j}}\in\mathbb{F}_{q}$ such that $(\sqrt{s_{j}})^{2}=s_{j}$. By this and (1.3) we obtain
(4.4) $a_{d}(q)=-2\lambda_{n/2}(d).$
Now combining (4) with (4.4), we obtain
$1\cdot 2^{n}\cdot\det T_{q}(d)=a_{d}(q)\cdot y_{q}^{2}.$
In view of the above, we have completed the proof.∎
Now we turn to the proof of our last result.
Proof of Corollary 1.1. Let notations be as above and let $p\equiv 1\ ({\rm{mod}}\ 4)$ be a prime with $p=2n+1$. When $p=5$, our result clearly holds.
Suppose now $p\geq 7$. As
$\lambda_{k}(d)=\frac{(-1)^{k}\chi_{p}^{-k}(d)}{2}\left(J_{p}(\phi,\chi_{q}^{k})-J_{p}(\phi,\chi_{p}^{-k})\right)$
for any $k\in[1,n-1]$, the rational number
$\frac{y_{p}}{2^{(n-2)/2}}={\bm{i}}^{(n-2)/2}\cdot\sqrt{p-1}\cdot\prod_{k\in(0,n/2)}\frac{1}{2}\left(J_{p}(\phi,\chi_{q}^{k})-J_{p}(\phi,\chi_{p}^{-k})\right)$
is indeed an algebraic integer. Thus,
(4.5) $z_{p}:=\frac{y_{p}}{2^{(n-2)/2}}\in\mathbb{Z}.$
On the other hand, for any $d\in\mathbb{F}_{p}^{\times}\setminus\mathcal{S}_{p}$, by (4.4) one can verify that
$(-1)^{n/2}\cdot J_{p}(\phi,\chi_{p}^{n/2})$ $=\sum_{j\in[1,n]}\phi(1+s_{j})\chi_{p}^{n/2}(s_{j})+\sum_{j\in[1,n]}\phi(1+ds_{j})\chi_{p}^{n/2}(ds_{j})$
$=\sum_{j\in[1,n]}\phi(1+s_{j})\phi(\sqrt{s_{j}})\pm{\bm{i}}\sum_{j\in[1,n]}\phi(1+ds_{j})\phi(\sqrt{s_{j}})$
$=\lambda_{n/2}(1)\pm{\bm{i}}\cdot\lambda_{n/2}(d)$
$=\lambda_{n/2}(1)\mp{\bm{i}}\cdot\frac{a_{p}(d)}{2}.$
Also, as $p\equiv 1\ ({\rm{mod}}\ 4)$, we see that
$\lambda_{n/2}(1)\equiv n-1\equiv 1\ ({\rm{mod}}\ 2).$
Hence, if we write $p=c_{p}^{2}+4b_{p}^{2}$ with $c_{p},b_{p}\in\mathbb{Z}^{+}$, then
(4.6) $\pm a_{p}(d)/4=b_{p}>0.$
Now applying Theorem 1.2, by (4.5) and (4.6) we see that
$\left|\det T_{d}(p)\right|/(2^{n}\cdot b_{p})=z_{p}^{2}$
is a square of the integer $z_{p}$. Moreover, Z.-W. Sun [11, Theorem 1.2(iii)] proved that
$\left(\frac{\det T_{d}(p)}{p}\right)=1$
whenever $d\in\mathbb{F}_{p}^{\times}\setminus\mathcal{S}_{p}$. This implies that $\det T_{d}(p)\in\mathbb{Z}\setminus\{0\}$.
In view of the above, $|\det T_{p}(d)|/(2^{(p-1)/2}b_{p})$ is indeed a positive integral square not depending on $d$. This completes the proof. ∎
5. Concluding Remarks
Let notations be as above. Recently, Z.-W. Sun [12, Conjecture 6.1] posed the following conjecture, which is a refinement of Conjecture 1.1.
Conjecture 5.1.
Let $p\equiv 1\ ({\rm{mod}}\ 4)$ be a prime. Then, there is an integer $t_{p}$ with $(\frac{t_{p}}{p})=1$ such that
$\det T_{p}(d)=2^{\frac{p-3}{2}}\cdot\left(\frac{p-1}{4}t_{p}\right)^{2}\cdot\sum_{j=1}^{\frac{p-1}{2}}\left(\frac{x^{3}+dx}{p}\right)$
for any $d\in\mathbb{F}_{p}^{\times}\setminus\mathcal{S}_{p}$.
In this section, we briefly discuss this stronger conjecture. Using (4.6), we first observe that
$2\sum_{j=1}^{(p-1)/2}\left(\frac{x^{2}+d}{p}\right)\left(\frac{x}{p}\right)$ $=\sum_{x\in\mathbb{F}_{p}}\left(\frac{x^{2}+d}{p}\right)\left(\frac{x}{p}\right)$
$=\sum_{x\in\mathbb{F}_{p}}\left(\frac{(dx)^{2}+d}{p}\right)\left(\frac{dx}{p}\right)$
$=\sum_{x\in\mathbb{F}_{p}}\left(\frac{dx^{2}+1}{p}\right)\left(\frac{x}{p}\right)$
$=-a_{p}(d)$
$=\pm 2b_{p}.$
Thus, Conjecture 5.1 is equivalent to saying that
$\frac{\left|\det T_{p}(d)\right|}{2^{(p-1)/2}\cdot b_{p}}=\left(\frac{p-1}{4}t_{p}\right)^{2}$
for some intger $t_{p}$ with $(\frac{t_{p}}{p})=1$. Moreover, by the proof of Corollary 1.1, we have proven that
$\frac{\left|\det T_{p}(d)\right|}{2^{(p-1)/2}\cdot b_{p}}=z_{p}^{2},$
where
$z_{p}=\frac{{\bm{i}}^{\frac{p-5}{4}}\cdot I_{p}(\chi_{p})\cdot\sqrt{p-1}}{2^{(p-5)/4}}\in\mathbb{Z}.$
In view of the above, one can verify that Conjecture 5.1 holds if and only if the rational number $4z_{p}/(p-1)$ is not only an integer but also a quadratic residue modulo $p$. However, we can not prove this result by the methods used in this paper.